# AURA: A Reinforcement Learning Framework for AI-Driven Adaptive Conversational Surveys

### v1

@@ Abstract
Conventional online surveys provide limited personalization, often resulting in low engagement and superficial responses. Although AI survey chatbots improve convenience, most are still reactive: they rely on fixed dialogue trees or static prompt templates and therefore cannot adapt within a session to fit individual users, which leads to generic follow-ups and weak response quality. We address these limitations with AURA (Adaptive Understanding through Reinforcement Learning for Assessment), a reinforcement learning framework for AI-driven adaptive conversational surveys. AURA quantifies response quality using a four-dimensional LSDE metric (Length, Self-disclosure, Emotion, and Specificity) and selects follow-up question types via an $\epsilon$-greedy policy that updates the expected quality gain (expected value) within each session. Initialized with priors extracted from 96 prior campus-climate conversations (467 total chatbot-user exchanges), the system balances exploration and exploitation across 10-15 dialogue exchanges, dynamically adapting to individual participants in real time. In controlled evaluations, AURA achieved a +0.12+0.12 mean gain in response quality and a statistically significant improvement over non-adaptive baselines ($p=0.044$, $d=0.66$), driven by a $63\%$ reduction in specification prompts and a $10\times$ increase in validation behavior. These results demonstrate that reinforcement learning can give survey chatbots improved adaptivity, transforming static questionnaires into interactive, self-improving assessment systems.
@@ 1 Introduction
Campus-climate surveys play a crucial role in understanding how students, faculty, and staff experience everyday life at universities. By gathering feedback on issues such as academic support, discrimination, resource accessibility, and sense of belonging, these surveys enable institutions to identify structural barriers and design interventions that promote equity and inclusion (Vogel et al., 2008; PASSHE, 2022; Bordieri et al., 2024). When administered on a regular cadence, they also provide a longitudinal view of institutional progress, ensuring that improvement efforts remain evidence based and responsive to community needs.
@@
However, the effectiveness of traditional survey instruments is constrained by their static format. Respondents often experience low engagement, survey fatigue, or limited opportunity for clarification, which can lead to superficial or incomplete data (Dillman et al., 2014; Galesic and Bosnjak, 2009; Sahlqvist et al., 2011). The impersonal, time consuming nature of conventional questionnaires further discourages participation and can undermine the reliability of collected information (Porter and Whitcomb, 2005). As a result, administrators may fail to capture the nuanced attitudes, emotions, and contextual factors that shape lived experiences on campus.
@@
Recent advances in generative artificial intelligence (GenAI) and large language models (LLMs) have created opportunities to transform these processes. Conversational AI systems can conduct interactive, context-aware dialogues that encourage deeper reflection (Xiao et al., 2020; Tanwar et al., 2024). Early deployments in campus settings demonstrate that chatbot-based surveys can yield richer qualitative feedback, improved user satisfaction, and reduced survey fatigue compared to traditional web forms (Abbas et al., 2021; Kim et al., 2019; Zarouali et al., 2024).
@@
Despite these advances, most existing AI-driven survey chatbots remain reactive rather than truly adaptive. Many still depend on rigid dialogue trees or static prompt templates that fails to evolve with user behavior during a session (Thorat and Jadhav, 2020; Maeng and Lee, 2021; Chen et al., 2024). Earlier frameworks such as TigerGPT (Tang et al., 2025), which combined role-based onboarding with empathetic conversational design, also followed pre-scripted flows. These systems showed that generative dialogue can boost engagement, yet they lacked a way to learn from user feedback and adjust questioning strategies as the conversation unfolded.
@@
This limitation narrows the potential of GenAI-driven surveys. Human interviewers naturally shift their style based on subtle cues — brevity, emotional tone, or how specific a response is — but static prompts cannot reproduce this kind of flexible reasoning. Work in dialogue management has shown that reinforcement learning can guide sequential decision making and improve both adaptability and personalization (Li et al., 2016; Su et al., 2021; Sun et al., 2023). More recent studies extend this approach to large language models, refining their decision policies through offline goal-conditioned optimization (Hong et al., 2025). Even so, reinforcement learning has rarely been applied to survey or assessment contexts, where adaptive behavior could meaningfully raise data quality and respondent trust.
@@
A persistent challenge is within-session personalization: how can a chatbot learn, inside a single exchange, which questioning strategy best fits each individual? Some respondents engage deeply when asked for concrete examples, while others prefer brief factual prompts. Emotional reflection may help some participants but discourage others. Current systems typically rely on fixed sequences or simple rule-based branching and therefore fail to recognize these individual patterns (Xiao et al., 2020; Belhaj et al., 2021). In campus-climate datasets, for instance, 64.2% of responses show low engagement (composite quality scores below 0.4), suggesting that static question logic struggles to sustain participation across diverse users.
@@
This gap highlights the missing connection between conversational design and computational learning. Human-computer-interaction frameworks emphasize empathy, personalization, and user-controlled flow (Kim et al., 2008; Kocaballi et al., 2019), whereas reinforcement learning formalizes decision making through reward feedback (Sutton et al., 1998). Bringing these perspectives together would allow an AI survey agent to align conversational empathy with measurable engagement outcomes — transforming reactive dialogue into an adaptive, continuously improving interaction process that better reflects human interviewing practice.
@@
To address this gap, we present AURA (Adaptive Understanding through Reinforcement Learning for Assessment), a reinforcement-learning–enhanced survey framework that learns optimal questioning strategies within individual conversations. AURA extends the conversational design principles of our pilot system, TigerGPT (Tang et al., 2025), with three key innovations:
@@
    Empirical Validation: Quantitative and behavioral analyses demonstrate a +0.12 mean gain in response quality and statistically significant improvements ($p=0.044$, $d=0.66$) under the optimal configuration ($\epsilon=0.30$, $\alpha=0.30$), confirming that within-session reinforcement learning enhances conversational-survey engagement.
@@ 2.1 Conversational Surveys and Chatbot-Based Data Collection
AI-powered chatbots are increasingly used in higher education for information gathering, feedback collection, and student engagement. Institutions deploy them to collect course evaluations, campus-climate assessments, and service-quality feedback, often achieving higher response rates and more detailed input than traditional web forms (Abbas et al., 2021; Belhaj et al., 2021). For example, pilots at European universities show that conversational survey interfaces improve participation and perceived personalization (Martinez-Requejo et al., 2024). These deployments suggest that, when designed around user needs, chatbots can transform routine data collection into an engaging, interactive experience.
@@
Within survey research, chat-based interfaces have emerged as an alternative to static questionnaires. Conversational surveys allow participants to respond in natural language and receive clarifying or follow-up questions in real time (Xiao et al., 2020; Tanwar et al., 2024). Studies consistently report higher engagement and richer qualitative responses compared with conventional forms (Abbas et al., 2021; Kim et al., 2019). For instance, Xiao et al. (2020) found that participants produced more nuanced and emotionally expressive feedback when interacting with a chatbot, while Tanwar et al. (2024) demonstrated that large language models (LLMs) can generate context-aware follow-up questions for class feedback. At the same time, not all users prefer automated dialogue. Zarouali et al. (2024) and Njeguš et al. (2021) observed that some respondents still favored conventional surveys because of privacy or trust concerns, highlighting sociocultural factors that influence the acceptance of AI-mediated data collection.
@@
The effectiveness of conversational surveys ultimately depends on the quality of the responses they elicit. Survey methodologists emphasize several indicators of response quality—length and elaboration (Dillman et al., 2014), emotional expression and engagement (Porter and Whitcomb, 2005), and the specificity of content (Brysbaert et al., 2014; Newman et al., 2003). Traditional questionnaires often fall short on these dimensions because they cannot adapt to respondent characteristics or probe for elaboration when needed. Chatbot interfaces offer the potential to improve response quality, but realizing this potential requires not only natural conversation but also strategic question selection that sustains engagement throughout the interaction.
@@
Early chatbot survey systems were typically rule-based, relying on predefined dialogue trees and keyword matching (Thorat and Jadhav, 2020; Maeng and Lee, 2021; Chan et al., 2022). Such rigid architectures offered little flexibility: unexpected inputs often led to conversational dead ends or irrelevant prompts. Recent advances in natural language processing and LLMs have enabled greater contextual understanding, allowing chatbots to interpret ambiguous language and maintain coherent multi-exchanges (Xiao et al., 2020; Tanwar et al., 2024). This shift marks a transition from deterministic scripts to generative, data-driven dialogue. Nevertheless, even current LLM-based survey systems remain largely reactive. They generate follow-ups based on local context but do not learn from user behavior within a session or across interactions. Furthermore, most dialogue research focuses on task-oriented conversations where success is defined by goal completion (e.g., booking reservations or retrieving information) (Su et al., 2021). In contrast, survey quality depends on sustained engagement and willingness to elaborate, which cannot be reduced to a binary success metric.
@@
This limitation is particularly evident in campus-climate surveys, where respondents vary widely in engagement patterns, disclosure comfort, and sensitivity to topics (Vogel et al., 2008; PASSHE, 2022). A first-year student discussing academic support may respond well to specific follow-up questions, whereas a faculty member describing discrimination may require validation and emotional acknowledgment before elaborating. Static prompts cannot accommodate such heterogeneity. While LLMs enable more natural language generation, they still lack a mechanism for learning which questioning strategies maintain engagement for each individual within a conversation. Addressing this limitation requires integrating conversational design with adaptive learning algorithms capable of optimizing question selection dynamically based on real-time feedback.
@@
Our prior work, TigerGPT (Tang et al., 2025), demonstrated that generative chatbots incorporating empathetic cues, bolded questions, and user-driven topic selection can enhance satisfaction and produce richer feedback than static surveys in campus-climate assessment. However, TigerGPT—and most comparable systems—operated with fixed conversational heuristics: questions followed predetermined templates based on user role (student, faculty, or staff) but did not adapt to engagement patterns observed during the dialogue. These observations motivate the development of AURA, a reinforcement learning framework that achieves within-session personalization through continuous quality feedback and adaptive action selection.
@@ 2.2 Reinforcement Learning for Dialogue Management
Reinforcement learning (RL) provides a principled framework for optimizing sequential decisions in interactive systems where explicit supervision is limited. In dialogue management, RL formulates conversation as a Markov Decision Process (MDP) in which an agent selects actions—responses, question types, or dialogue strategies—based on conversational states to maximize cumulative reward such as task success or user satisfaction (Sutton et al., 1998). This paradigm enables systems to improve their policies through experience rather than relying solely on predefined rules.
Early research applied RL to task-oriented dialogues such as restaurant reservations or technical support. Li et al. (2016) pioneered deep RL for open-domain conversation, showing that policy-gradient optimization could enhance coherence and diversity beyond supervised models. Su et al. (2021) extended this framework through dynamic policy networks that adapt questioning strategies in real time for slot-filling tasks. Subsequent studies introduced contextual bandits to improve personalization and sample efficiency (Sun et al., 2023), while recent work integrates RL with large language models via offline goal-conditioned learning to refine model policies without costly online trial-and-error (Hong et al., 2025). Together, these advances illustrate RL’s potential for adaptive dialogue.
Within-session learning. Most RL dialogue systems employ multi-session learning: a global policy improves as the agent interacts with many users over time (Li et al., 2016; Su et al., 2021). Survey applications violate this assumption—each respondent interacts only once, exploration is costly, and privacy constraints limit cross-user adaptation (Kocaballi et al., 2019). Within-session learning instead focuses on rapid adaptation inside a single conversation. The agent begins from prior knowledge derived from prior data and adjusts its strategy in real time based on user engagement, prioritizing individual responsiveness over long-term population learning. Algorithms that converge within 10–15 exchanges are therefore required, motivating lightweight tabular or expected-value methods rather than deep networks.
@@ 2.3 Quality Assessment in Open-Ended Surveys
Prior research in survey methodology emphasizes that open-ended response quality is multidimensional rather than captured by a single indicator. Empirical and psycholinguistic studies identify four recurring factors: response length, self-disclosure, emotional expression, and specificity. Together, these features shape engagement and informativeness in qualitative data (Dillman et al., 2014; Holland and Christian, 2009; Pennebaker et al., 2003; Tausczik and Pennebaker, 2010; Brysbaert et al., 2014). These dimensions have been used in both human-coded (Krosnick, 1991; Holbrook et al., 2003) and automated text-analytic assessments. Recent conversational-survey systems similarly exploit lexical and emotional cues to estimate engagement in real time (Xiao et al., 2020; Tanwar et al., 2024), motivating integration of such dynamic linguistic feedback within adaptive questioning frameworks. This body of work forms the theoretical foundation for AURA’s linguistic quality metric introduced in Section 3.3, which operationalizes these dimensions for real-time evaluation within conversational surveys.
@@ 2.4 Individual Differences and the Need for Adaptive Questioning
Heterogeneity in survey response patterns. Research in survey methodology consistently demonstrates diverse engagement styles in open-ended responses. Krosnick (1991) identified systematic differences in response elaboration—some individuals naturally produce detailed narratives, while others provide brief replies even when equally knowledgeable. Holbrook et al. (2003) found that question-format effectiveness varies by respondent characteristics, with certain demographic groups responding better to specific than to general prompts. Tourangeau et al. (2000) showed that cognitive load, which fluctuates both across individuals and within conversations, affects how people process questions and formulate answers. Together, these findings suggest that optimal questioning complexity must adapt to each respondent’s cognitive resources at any given moment.
@@
Cognitive and affective factors. Several psychological mechanisms underlie these variations. Cognitive load theory (Lenzner et al., 2010) posits that question complexity interacts with working-memory capacity: respondents with lower available resources benefit from simpler, structured prompts, whereas those with higher capacity may disengage from overly basic questions. Affective factors further shape engagement. Melville and Hincks (2016) found that trust and willingness to self-disclose differ substantially across individuals and evolve throughout sensitive interviews. Some participants require rapport building before sharing personal experiences, while others engage immediately. Wheeless and Grotz (1977) demonstrated that disclosure patterns are both person- and context-dependent, implying that effective interviewers must continuously calibrate their approach based on real-time feedback cues.
@@
Communication styles and individual preferences. Individual differences also appear in preferred communication styles. Gudykunst and Ting-Toomey (1988) showed that people vary in receptiveness to direct versus indirect questioning, with some favoring explicit requests for information and others responding better to open-ended exploration. Preferences for elaborative versus succinct exchanges likewise differ: some respondents welcome follow-up probes that expand on previous answers, while others perceive repeated questioning as intrusive or redundant (Kim et al., 2019). Characteristics such as communication style, cultural background, educational experience, and topic familiarity all influence engagement patterns (Pennebaker and King, 1999; Gudykunst and Ting-Toomey, 1988), yet the specific combination varies by individual and context. Such variability indicates that effective conversational systems cannot rely on demographic stereotypes or population-level averages—they must infer user preferences dynamically during the interaction.
@@
Evidence from conversational systems. Similar patterns appear in chatbot-based surveys. Zarouali et al. (2024) reported that response quality varied more by individual than by question type, underscoring the limitations of uniform dialogue strategies. In healthcare contexts, Kocaballi et al. (2019) showed that personalized conversational agents achieve higher satisfaction and greater information disclosure than generic systems, though most rely on explicit preference elicitation rather than real-time behavioral adaptation.
@@
Limitations of one-size-fits-all approaches. Our pilot analysis provides direct evidence of these individual differences. The TigerGPT system (Tang et al., 2025) employed consistent conversational strategies across 96 campus-climate conversations, following empathetic design principles and user-driven topic selection. Yet among 467 responses collected, 64.2% exhibited low engagement (composite quality scores below 0.4). Response quality varied substantially across users: scores ranged from 0.0 to 1.0, with some respondents providing rich detail (above 0.6, 16.7 %) and others giving minimal one-word answers (below 0.2, 47.1 %). This heterogeneity demonstrates that even well-designed conversational systems with empathetic cues cannot succeed using fixed strategies. Effective adaptation requires learning individual preferences in real time from engagement patterns, motivating AURA’s adaptive learning framework for discovering optimal questioning strategies within each single conversation.

### v2

@@ Abstract
Conventional online surveys provide limited personalization, often resulting in low engagement and superficial responses. Although AI survey chatbots improve convenience, most are still reactive: they rely on fixed dialogue trees or static prompt templates and therefore cannot adapt within a session to fit individual users, which leads to generic follow-ups and weak response quality. We address these limitations with AURA (Adaptive Understanding through Reinforcement Learning for Assessment), a reinforcement learning framework for AI-driven adaptive conversational surveys. AURA quantifies response quality using a four-dimensional LSDE metric (Length, Self-disclosure, Emotion, and Specificity) and selects follow-up question types via an $\epsilon$-greedy policy that updates the expected quality gain (expected value) within each session. Initialized with priors extracted from 96 prior campus-climate conversations (467 total chatbot-user exchanges), the system balances exploration and exploitation across 10-15 dialogue exchanges, dynamically adapting to individual participants in real time. In controlled evaluations, AURA achieved a +0.076+0.076 mean gain in response quality and a statistically significant improvement over non-adaptive baselines ($p=0.044$, $d=0.66$), driven by a $63\%$ reduction in specification prompts and a $10\times$ increase in validation behavior. These results demonstrate that reinforcement learning can give survey chatbots improved adaptivity, transforming static questionnaires into interactive, self-improving assessment systems.
@@ 1 Introduction
Campus-climate surveys play a crucial role in understanding how students, faculty, and staff experience everyday life at universities. By gathering feedback on issues such as academic support, discrimination, resource accessibility, and sense of belonging, these surveys enable institutions to identify structural barriers and design interventions that promote equity and inclusion (vogel2008assessment; PASSHE2022; bordieri2024exploring). When administered on a regular cadence, they also provide a longitudinal view of institutional progress, ensuring that improvement efforts remain evidence based and responsive to community needs.
@@
However, the effectiveness of traditional survey instruments is constrained by their static format. Respondents often experience low engagement, survey fatigue, or limited opportunity for clarification, which can lead to superficial or incomplete data (dillman2014internet; galesic2009effects; sahlqvist2011effect). The impersonal, time consuming nature of conventional questionnaires further discourages participation and can undermine the reliability of collected information (porter2005mail). As a result, administrators may fail to capture the nuanced attitudes, emotions, and contextual factors that shape lived experiences on campus.
@@
Recent advances in generative artificial intelligence (GenAI) and large language models (LLMs) have created opportunities to transform these processes. Conversational AI systems can conduct interactive, context-aware dialogues that encourage deeper reflection (xiao2020tell; tanwar2024opinebot). Early deployments in campus settings demonstrate that chatbot-based surveys can yield richer qualitative feedback, improved user satisfaction, and reduced survey fatigue compared to traditional web forms (abbas2021university; kim2019comparing; zarouali2024comparing).
@@
Despite these advances, most existing AI-driven survey chatbots remain reactive rather than truly adaptive. Many still depend on rigid dialogue trees or static prompt templates that fails to evolve with user behavior during a session (thorat2020review; maeng2021designing; chen2024recent). Earlier frameworks such as TigerGPT (tang2025tigergpt), which combined role-based onboarding with empathetic conversational design, also followed pre-scripted flows. These systems showed that generative dialogue can boost engagement, yet they lacked a way to learn from user feedback and adjust questioning strategies as the conversation unfolded.
@@
This limitation narrows the potential of GenAI-driven surveys. Human interviewers naturally shift their style based on subtle cues — brevity, emotional tone, or how specific a response is — but static prompts cannot reproduce this kind of flexible reasoning. Work in dialogue management has shown that reinforcement learning can guide sequential decision making and improve both adaptability and personalization (li2016deep; su2021dynamic; sun2023contextual). More recent studies extend this approach to large language models, refining their decision policies through offline goal-conditioned optimization (hong2025planning). Even so, reinforcement learning has rarely been applied to survey or assessment contexts, where adaptive behavior could meaningfully raise data quality and respondent trust.
@@
A persistent challenge is within-session personalization: how can a chatbot learn, inside a single exchange, which questioning strategy best fits each individual? Some respondents engage deeply when asked for concrete examples, while others prefer brief factual prompts. Emotional reflection may help some participants but discourage others. Current systems typically rely on fixed sequences or simple rule-based branching and therefore fail to recognize these individual patterns (xiao2020tell; belhaj2021engaging). In campus-climate datasets, for instance, 64.2% of responses show low engagement (composite quality scores below 0.4), suggesting that static question logic struggles to sustain participation across diverse users.
@@
This gap highlights the missing connection between conversational design and computational learning. Human-computer-interaction frameworks emphasize empathy, personalization, and user-controlled flow (kim2008keeping; kocaballi2019personalization), whereas reinforcement learning formalizes decision making through reward feedback (sutton1998reinforcement). Bringing these perspectives together would allow an AI survey agent to align conversational empathy with measurable engagement outcomes — transforming reactive dialogue into an adaptive, continuously improving interaction process that better reflects human interviewing practice.
@@
To address this gap, we present AURA (Adaptive Understanding through Reinforcement Learning for Assessment), a reinforcement-learning–enhanced survey framework that learns optimal questioning strategies within individual conversations. AURA extends the conversational design principles of our pilot system, TigerGPT (tang2025tigergpt), with three key innovations:
@@
    Empirical Validation: Quantitative and behavioral analyses demonstrate a +0.076 mean gain in response quality and statistically significant improvements ($p=0.044$, $d=0.66$) under the optimal configuration ($\epsilon=0.30$, $\alpha=0.30$), confirming that within-session reinforcement learning enhances conversational-survey engagement.
@@ 2.1 Conversational Surveys and Chatbot-Based Data Collection
AI-powered chatbots are increasingly used in higher education for information gathering, feedback collection, and student engagement. Institutions deploy them to collect course evaluations, campus-climate assessments, and service-quality feedback, often achieving higher response rates and more detailed input than traditional web forms (abbas2021university; belhaj2021engaging). For example, pilots at European universities show that conversational survey interfaces improve participation and perceived personalization (martinez2024ai). These deployments suggest that, when designed around user needs, chatbots can transform routine data collection into an engaging, interactive experience.
@@
Within survey research, chat-based interfaces have emerged as an alternative to static questionnaires. Conversational surveys allow participants to respond in natural language and receive clarifying or follow-up questions in real time (xiao2020tell; tanwar2024opinebot). Studies consistently report higher engagement and richer qualitative responses compared with conventional forms (abbas2021university; kim2019comparing). For instance, xiao2020tell found that participants produced more nuanced and emotionally expressive feedback when interacting with a chatbot, while tanwar2024opinebot demonstrated that large language models (LLMs) can generate context-aware follow-up questions for class feedback. At the same time, not all users prefer automated dialogue. zarouali2024comparing and njeguvs2021conversational observed that some respondents still favored conventional surveys because of privacy or trust concerns, highlighting sociocultural factors that influence the acceptance of AI-mediated data collection.
@@
The effectiveness of conversational surveys ultimately depends on the quality of the responses they elicit. Survey methodologists emphasize several indicators of response quality—length and elaboration (dillman2014internet), emotional expression and engagement (porter2005mail), and the specificity of content (brysbaert2014concreteness; newman2003lying). Traditional questionnaires often fall short on these dimensions because they cannot adapt to respondent characteristics or probe for elaboration when needed. Chatbot interfaces offer the potential to improve response quality, but realizing this potential requires not only natural conversation but also strategic question selection that sustains engagement throughout the interaction.
@@
Early chatbot survey systems were typically rule-based, relying on predefined dialogue trees and keyword matching (thorat2020review; maeng2021designing; chan2022challenges). Such rigid architectures offered little flexibility: unexpected inputs often led to conversational dead ends or irrelevant prompts. Recent advances in natural language processing and LLMs have enabled greater contextual understanding, allowing chatbots to interpret ambiguous language and maintain coherent multi-exchanges (xiao2020tell; tanwar2024opinebot). This shift marks a transition from deterministic scripts to generative, data-driven dialogue. Nevertheless, even current LLM-based survey systems remain largely reactive. They generate follow-ups based on local context but do not learn from user behavior within a session or across interactions. Furthermore, most dialogue research focuses on task-oriented conversations where success is defined by goal completion (e.g., booking reservations or retrieving information) (su2021dynamic). In contrast, survey quality depends on sustained engagement and willingness to elaborate, which cannot be reduced to a binary success metric.
@@
This limitation is particularly evident in campus-climate surveys, where respondents vary widely in engagement patterns, disclosure comfort, and sensitivity to topics (vogel2008assessment; PASSHE2022). A first-year student discussing academic support may respond well to specific follow-up questions, whereas a faculty member describing discrimination may require validation and emotional acknowledgment before elaborating. Static prompts cannot accommodate such heterogeneity. While LLMs enable more natural language generation, they still lack a mechanism for learning which questioning strategies maintain engagement for each individual within a conversation. Addressing this limitation requires integrating conversational design with adaptive learning algorithms capable of optimizing question selection dynamically based on real-time feedback.
@@
Our prior work, TigerGPT (tang2025tigergpt), demonstrated that generative chatbots incorporating empathetic cues, bolded questions, and user-driven topic selection can enhance satisfaction and produce richer feedback than static surveys in campus-climate assessment. However, TigerGPT—and most comparable systems—operated with fixed conversational heuristics: questions followed predetermined templates based on user role (student, faculty, or staff) but did not adapt to engagement patterns observed during the dialogue. These observations motivate the development of AURA, a reinforcement learning framework that achieves within-session personalization through continuous quality feedback and adaptive action selection.
@@ 2.2 Reinforcement Learning for Dialogue Management
Reinforcement learning (RL) provides a principled framework for optimizing sequential decisions in interactive systems where explicit supervision is limited. In dialogue management, RL formulates conversation as a Markov Decision Process (MDP) in which an agent selects actions—responses, question types, or dialogue strategies—based on conversational states to maximize cumulative reward such as task success or user satisfaction (sutton1998reinforcement). This paradigm enables systems to improve their policies through experience rather than relying solely on predefined rules.
Early research applied RL to task-oriented dialogues such as restaurant reservations or technical support. li2016deep pioneered deep RL for open-domain conversation, showing that policy-gradient optimization could enhance coherence and diversity beyond supervised models. su2021dynamic extended this framework through dynamic policy networks that adapt questioning strategies in real time for slot-filling tasks. Subsequent studies introduced contextual bandits to improve personalization and sample efficiency (sun2023contextual), while recent work integrates RL with large language models via offline goal-conditioned learning to refine model policies without costly online trial-and-error (hong2025planning). Together, these advances illustrate RL’s potential for adaptive dialogue.
Within-session learning. Most RL dialogue systems employ multi-session learning: a global policy improves as the agent interacts with many users over time (li2016deep; su2021dynamic). Survey applications violate this assumption—each respondent interacts only once, exploration is costly, and privacy constraints limit cross-user adaptation (kocaballi2019personalization). Within-session learning instead focuses on rapid adaptation inside a single conversation. The agent begins from prior knowledge derived from prior data and adjusts its strategy in real time based on user engagement, prioritizing individual responsiveness over long-term population learning. Algorithms that converge within 10–15 exchanges are therefore required, motivating lightweight tabular or expected-value methods rather than deep networks.
@@ 2.3 Quality Assessment in Open-Ended Surveys
Prior research in survey methodology emphasizes that open-ended response quality is multidimensional rather than captured by a single indicator. Empirical and psycholinguistic studies identify four recurring factors: response length, self-disclosure, emotional expression, and specificity. Together, these features shape engagement and informativeness in qualitative data (dillman2014internet; holland2009measuring; pennebaker2003linguistic; tausczik2010psychological; brysbaert2014concreteness). These dimensions have been used in both human-coded (krosnick1991response; holbrook2003telephone) and automated text-analytic assessments. Recent conversational-survey systems similarly exploit lexical and emotional cues to estimate engagement in real time (xiao2020tell; tanwar2024opinebot), motivating integration of such dynamic linguistic feedback within adaptive questioning frameworks. This body of work forms the theoretical foundation for AURA’s linguistic quality metric introduced in Section 3.3, which operationalizes these dimensions for real-time evaluation within conversational surveys.
@@ 2.4 Individual Differences and the Need for Adaptive Questioning
Heterogeneity in survey response patterns. Research in survey methodology consistently demonstrates diverse engagement styles in open-ended responses. krosnick1991response identified systematic differences in response elaboration—some individuals naturally produce detailed narratives, while others provide brief replies even when equally knowledgeable. holbrook2003telephone found that question-format effectiveness varies by respondent characteristics, with certain demographic groups responding better to specific than to general prompts. tourangeau2000psychology showed that cognitive load, which fluctuates both across individuals and within conversations, affects how people process questions and formulate answers. Together, these findings suggest that optimal questioning complexity must adapt to each respondent’s cognitive resources at any given moment.
@@
Cognitive and affective factors. Several psychological mechanisms underlie these variations. Cognitive load theory (lenzner2010cognitive) posits that question complexity interacts with working-memory capacity: respondents with lower available resources benefit from simpler, structured prompts, whereas those with higher capacity may disengage from overly basic questions. Affective factors further shape engagement. melville2016conducting found that trust and willingness to self-disclose differ substantially across individuals and evolve throughout sensitive interviews. Some participants require rapport building before sharing personal experiences, while others engage immediately. wheeless1977measurement demonstrated that disclosure patterns are both person- and context-dependent, implying that effective interviewers must continuously calibrate their approach based on real-time feedback cues.
@@
Communication styles and individual preferences. Individual differences also appear in preferred communication styles. gudykunst1988culture showed that people vary in receptiveness to direct versus indirect questioning, with some favoring explicit requests for information and others responding better to open-ended exploration. Preferences for elaborative versus succinct exchanges likewise differ: some respondents welcome follow-up probes that expand on previous answers, while others perceive repeated questioning as intrusive or redundant (kim2019comparing). Characteristics such as communication style, cultural background, educational experience, and topic familiarity all influence engagement patterns (pennebaker1999linguistic; gudykunst1988culture), yet the specific combination varies by individual and context. Such variability indicates that effective conversational systems cannot rely on demographic stereotypes or population-level averages—they must infer user preferences dynamically during the interaction.
@@
Evidence from conversational systems. Similar patterns appear in chatbot-based surveys. zarouali2024comparing reported that response quality varied more by individual than by question type, underscoring the limitations of uniform dialogue strategies. In healthcare contexts, kocaballi2019personalization showed that personalized conversational agents achieve higher satisfaction and greater information disclosure than generic systems, though most rely on explicit preference elicitation rather than real-time behavioral adaptation.
@@
Limitations of one-size-fits-all approaches. Our pilot analysis provides direct evidence of these individual differences. The TigerGPT system (tang2025tigergpt) employed consistent conversational strategies across 96 campus-climate conversations, following empathetic design principles and user-driven topic selection. Yet among 467 responses collected, 64.2% exhibited low engagement (composite quality scores below 0.4). Response quality varied substantially across users: scores ranged from 0.0 to 1.0, with some respondents providing rich detail (above 0.6, 16.7 %) and others giving minimal one-word answers (below 0.2, 47.1 %). This heterogeneity demonstrates that even well-designed conversational systems with empathetic cues cannot succeed using fixed strategies. Effective adaptation requires learning individual preferences in real time from engagement patterns, motivating AURA’s adaptive learning framework for discovering optimal questioning strategies within each single conversation.